OpenStudy (anonymous):

the sum of the digits in a two digit number is 12, if the digits are reversed, the new number is 18 more than the original, find the original number

jimthompson5910 (jim_thompson5910):

Let t = tens digit u = units digit we are told that "the sum of the digits in a two digit number is 12" so this means t + u = 12

jimthompson5910 (jim_thompson5910):

Solve for t to get t + u = 12 t = 12 - u

jimthompson5910 (jim_thompson5910):

Then we're told that "if the digits are reversed, the new number is 18 more than the original" which translates to this equation 10u+t = 10t+u + 18 Now plug in t = 12 - u to get 10u+12-u = 10(12-u)+u + 18 From here, you solve for u. Tell me what you get.

OpenStudy (anonymous):

im stupid so I cant solve that

jimthompson5910 (jim_thompson5910):

I'll get you started 10u+12-u = 10(12-u)+u + 18 10u+12-u = 120-10u+u + 18 9u + 12 = -9u + 138 u = ???

OpenStudy (anonymous):

is the original number 75?

jimthompson5910 (jim_thompson5910):

close, it's actually 57

OpenStudy (anonymous):

thank you

jimthompson5910 (jim_thompson5910):

you're welcome